# NEST: Self-supervised Fast Conformer as All-purpose Seasoning to Speech Processing Tasks

### v5

@@ I Introduction
HuBERT [6], as a pioneer work of the predictive approach, generates the target tokens by running k-means clustering on the middle layer features extracted from anohter SSL model that is pretrained for a small number of steps. Then, W2v-BERT [14] proposes to combine the training objectives of both Wav2vec-2.0 [2] and HuBERT [6] by applying contrastive loss on the middle layer output and predictive loss on the final output layer. Later, BEST-RQ [7] shows that the clustering based token generation can be replaced by simple fixed random-projection quantization, and this simple modification is able to match or outperform HuBERT on ASR.
@@
  • Experiment results demonstrate that NEST can help achieve SOTA performance on a variety of downstream tasks (ASR, AST, SLU, SD, etc).
  • Unlike previous SSL approaches that primarily focus on downstream tasks with limited data, we also show that NEST can benefit speech recognition and translation even when data is relatively larger.
@@ II-C Speech Quantization
We use BEST-RQ [7] for speech quantization. Specifically, we employ a single randomly initialized and frozen codebook of 8192 vocabulary and 16 dimension features. A randomly initialized and frozen linear layer is applied to the input Mel-spectrogram features to project them into the same dimension as the codebook, then a nearest neighbor search is applied to obtain the target tokens. Since there is a 8x subsampling, we channel-concatenate the features for each 8 consecutive frames before feeding into the linear layer, such that the lengths for the target tokens and input features are equal.
@@ III-A Dataset and Settings
We train the NEST-L (108M) and NEST-XL (600M) models using 100K hours of English speech data, including 60K hours from LibriLight [22], 24K hours from English subset of Voxpopuli [23], and about 20K hours sampled data from the combination of Fisher [24], Switchboard [25], WSJ [26], NSC [27], People’s Speech [28]. The audios for speech augmentation are randomly selected within each batch, while we use noise audios from MUSAN [29] and Freesound [30]. We train the models with global batch size of 2048 for 80K steps on 128 NVIDIA A100 GPUs, with Noam annealing [16] and peak learning rate of 0.004, weight decay of 1e-3, gradient clipping 1.0 and warm-up of 25K steps. We set the speech augmentation probability as 0.2, among which we set noise and speech augmentation probabilities as 0.1 and 0.9.
@@ III-B Results on SUPERB Multi-task Speech Processing
NEST-L-Sortformer-HL [37] 16.06 6.40 11.02 16.85
NEST-L-Sortformer-HL-PP [37] 14.60 6.08 9.57 15.40
@@ III-C Results on Multi-lingual Speech Recognition
From the last two rows of Table II, we can see that NEST can help achieve better WER on all datasets than the model with ASR pretrained initialization, which shows that NEST can help improve ASR performance on languages that is not seen during SSL pretraining. In addition, when compared with other SOTA ASR models (Whisper [20], SeamlessM4T [19], Canary [21]) trained with much more parameters and data, we are still able to match the performance of Canary [21] on averaged WER across all languages. On some of the datasets, although there is still a gap between our model’s performance and that of the SOTA models trained with much more data, we can still see that NEST can be used as an efficient way to obtain superior ASR performance comparable to models trained on massive datasets.

### v6

@@ I Introduction
HuBERT [6], as a pioneer work of the predictive approach, generates the target tokens by running k-means clustering on the middle layer features extracted from another SSL model that is pretrained for a small number of steps. Then, W2v-BERT [14] proposes to combine the training objectives of both Wav2vec-2.0 [2] and HuBERT [6] by applying contrastive loss on the middle layer output and predictive loss on the final output layer. Later, BEST-RQ [7] shows that the clustering based token generation can be replaced by simple fixed random-projection quantization, and this simple modification is able to match or outperform HuBERT on ASR.
@@
  • Experiment results demonstrate that NEST can help achieve SOTA performance on a variety of downstream tasks (ASR, AST, SD, SLU etc).
  • Unlike previous SSL approaches that primarily focus on downstream tasks with very limited data, we also show that NEST can benefit speech recognition and translation even when data is relatively larger.
@@ II-C Speech Quantization
We use BEST-RQ [7] for speech quantization. Specifically, we employ a single randomly initialized and frozen codebook of 8192 vocabulary and 16 dimension features. A randomly initialized and frozen linear layer is applied to the input Mel-spectrogram features to project them into the same dimension as the codebook, then a nearest neighbor search is applied to obtain the target tokens. Since there is an 8x subsampling, we channel-concatenate the features for each 8 consecutive frames before feeding into the linear layer, such that the lengths for the target tokens and input features are equal.
@@ III-A Dataset and Settings
We train the NEST-L (115M) and NEST-XL (600M) models using 100K hours of English speech data, including 60K hours from LibriLight [22], 24K hours from English subset of Voxpopuli [23], and about 20K hours sampled data from the combination of Fisher [24], Switchboard [25], WSJ [26], NSC [27], People’s Speech [28]. The audios for speech augmentation are randomly selected within each batch for each sample, and we use non-vocal noise audios from MUSAN [29] and Freesound [30]. We train the models with global batch size of 2048 samples for about 800K steps on 128 NVIDIA A100 GPUs, with Noam annealing [16], peak learning rate of 0.004, weight decay of 1e-3, gradient clipping 1.0 and warm-up of 25K steps. We set the speech augmentation probability as 0.2, among which we set noise and speech augmentation probabilities as 0.1 and 0.9.
@@ III-B Results on SUPERB Multi-task Speech Processing
NEST-L-Sortformer-HL [37] 16.28 6.49 10.01 14.14
NEST-L-Sortformer-HL-PP [37] 14.76 5.87 8.46 12.59
@@ III-C Results on Multi-lingual Speech Recognition
From the last two rows of Table II, we can see that NEST can help achieve better WER on all datasets than the model with ASR pretrained initialization, which shows that NEST can help improve ASR performance on languages that is not seen during SSL pretraining. In addition, when compared with other SOTA ASR models (Whisper [20], SeamlessM4T [19], Canary [21]) trained with much more parameters and data, we are still able to match the performance of Canary [21] on averaged WER across all languages. On some of the datasets, although there is still a gap between our model’s performance and that of the SOTA models trained with much more data, we can still see that NEST can be used as an efficient way to obtain good ASR performance comparable to models trained on massive datasets.